The points which trisect the line segment joining the points (0,0) and (9,12) are

Grade:8

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Please find the link below

https://www.askiitians.com/forums/Analytical-Geometry/24/7486/rc-2.htm

Regards
Vikas TU
4 years ago
Suppose Pint P and Q trisect the line segment joinng the points (0.0) and (9,12) 
P divides ABin ratio 1:2 and Q divides in ratio 2:1.
Point (x,y) divides the line segment joining the points (x1,y1) and (x2,y2) internally in the ratio m:n then 
(x,y) = (mx2 + nx1)/(m+n) + (my2+ny1)/(m+n) 
substituting m = 1 and n =2 (x,y) = (3,4)
substituting m = 2 and n =1 (x,y) = (6,8)